OpenStudy (anonymous):

What are good x values to choose for graphing y = 0.5x^2 ?

OpenStudy (anonymous):

You may pick any x value that is most convenient for you. Ideally, small numbers. ^_^ You don't have to worry about finding GOOD x-values; the equation will sort out the information itself. You just find what y equals when you give x a value, and you'll see the graph come to life. :D

OpenStudy (anonymous):

Lol, how about -1, -2, 0, 1, 2?

OpenStudy (anonymous):

Yeah. All o' that works.

OpenStudy (ranga):

How about [-6, +6] in steps of 1?

OpenStudy (ranga):

Since x is squared, the graph will be symmetric with respect to the y-axis. So x = -5 and x = +5 will give the same y value.

OpenStudy (anonymous):

I need at 5 x values. I need to make a quadratic graph.

OpenStudy (anonymous):

Ingore the at

OpenStudy (ranga):

Your choice of -1, -2, 0, 1, 2 will be fine. But you don't have to calculate for negative x values because x = -2 and x = +2 will give the same y value. So just calculate on the positive side and just put the same value on the corresponding negative x.

OpenStudy (anonymous):

Wouldn't it make it easier if I did that?

OpenStudy (anonymous):

Or by doing that I should say.

OpenStudy (ranga):

Just go ahead and plug in x = -1, -2, 0, 1, 2 and calculate the y values and plot them.

OpenStudy (anonymous):

Will you check my graph to see if I'm correct?

OpenStudy (ranga):

I have to leave in a few minutes. But if you will post it soon I will take a look.

OpenStudy (anonymous):

Ill try to hurry!

OpenStudy (ranga):

Based on the equation, do you know the name for this graph?

OpenStudy (anonymous):

A quadratic graph no?

OpenStudy (ranga):

Yes, it is a quadratic graph but the shape is called a parabola.

OpenStudy (anonymous):

\[.5(-2)^2=2\]

OpenStudy (ranga):

yes.

OpenStudy (anonymous):

\[.5(-1)^2=.5\]

OpenStudy (ranga):

correct. go ahead and type the rest of it in one reply.

OpenStudy (anonymous):

\[.5(0)^2=0\] \[.5(1)^2=.5\] \[.5(2)^2=2\]

OpenStudy (ranga):

Correct. So what are the x and y coordinates of the 5 points that you are going to plot?

OpenStudy (anonymous):

(-2,2) (-1,.5) (0,0) (1,.5) (2,2)

OpenStudy (ranga):

correct.

OpenStudy (anonymous):

I have to draw the graph now, let me try my best

OpenStudy (anonymous):